# Nearly optimal spectral gaps for random Belyi surfaces

### v1

@@ 2.3. An alternative description of Brooks-Makover model
For $n\geq 1$, assume $S_{6n}$ is the permutation group of $6n$ elements. Now we consider a special subset of $S_{6n}\times S_{6n}$, which is defined as
@@
    of $T_{k}$. Then glue the sides $j_{s}^{1}$ and $j_{s}^{2}\ (1\leq s\leq 3n)$ together subject to precise two conditions;
@@
for any $\gamma\in\textnormal{PSL}(2,\mathbb{Z}),\ x\in\mathbb{H}$ and $i\in[6n]$.
@@ 3.2.1. Injective $X$-labeled graph homomorphisms
The following lemma will be applied later.
@@
(([MPvH25, Lemma 2.2])) Let P,QP,\ Q be real polynomials of degree at most $m\in\mathbb{N}$, and let
@@
for all $n\in\mathbb{N}$ with $n\geq(Cm)^{C}$. Then there is a constant $C^{\prime}=C^{\prime}(C)>1$ such that for all $l\leq Cm$,
@@
    There exists a universal constant $C_{0}>1$ such that
    $\left|\mathbb{E}_{n}(N_{\Gamma})-\frac{1}{(6n)^{\eta(\Gamma)}}\sum\limits_{i=0}^{l-1}\frac{a_{i}(\Gamma)}{(6n)^{i}}\right|\leq\frac{(C_{0}m)^{C_{0}l}}{l!(6n)^{l+\eta(\Gamma)}}$
    for all $m\geq 2k$, $l\leq m$ and $6n\geq(C_{0}m)^{C_{0}}$.
@@
(1). Assume $\Gamma$ has $p$ $x_{1}$-cycles and $q$ $x_{2}$-cycles. Since $(g,\Gamma)\in\mathcal{I}(\omega)$, we have
@@
(2). Assume $m\geq 2k\geq 2p$. For
$C>3>\left(\prod\limits_{i=2}^{\infty}\left(1-\frac{1}{i^{2}}\right)\right)^{-1}$
and $6n\geq(Cm)^{C}$, we have
$\left|P\left(\frac{1}{6n}\right)\right|<1\leq C$
@@
$C>1$ $>Q\left(\frac{1}{6n}\right)=\prod\limits_{i=0}^{2p-1}\left(1-\frac{i}{6n}\right)$
$\geq\prod\limits_{i=0}^{2p-1}\left(1-\frac{i}{(2pC)^{C}}\right)\geq\prod\limits_{i=0}^{2p-1}\left(1-\frac{1}{i^{2}}\right)\geq\frac{1}{C}.$
Apply Lemma 3.3 to P(t),Q(t)P(t),\ Q(t) and $C$, there exists a constant $C_{0}=C_{0}(C)>1$ such that for any $l\leq Cm$,
$\sup\limits_{t\in\left[0,(C_{0}m)^{-C_{0}}\right]}\left|\left(\frac{P}{Q}\right)^{(l)}(t)\right|\leq(C_{0}m)^{C_{0}l}.$
It follows that for $6n\geq(C_{0}m)^{C_{0}}$ and $l\leq m$, there exists $\xi\in[0,\frac{1}{6n}]$ such that
@@ 3.2.2. Surjective $X$-labeled graph homomorphisms
Recall from the algorithm above, the initial data $\mathcal{A}_{1}=(\Gamma_{1},G_{1},f_{1},a_{1})$, and after the $t$-th step ($1\leq t\leq k$), we obtain
@@
Since there are at most $\binom{k}{h}$ different choices to determine all the $h$ steps at which we make Operation II, it follows that
$|\mathfrak{F}_{h}(\omega)|\leq\binom{k}{h}\times k^{h}\leq k^{2h},$
@@ 3.2.3. Number of fixed points
where $k\geq 1$ and $i_{j}\in\{1,2\}\ (1\leq j\leq k)$. Then there exist a universal constant $C\geq 1$ and a sequence $\{v_{i}(\omega)\}_{i\geq 0}\subset\mathbb{R}$ such that
$\left|\mathbb{E}_{n}\left[\textnormal{fix}_{\omega}\right]-\sum\limits_{i=0}^{m-1}\frac{v_{i}(\omega)}{(6n)^{i}}\right|\leq\frac{(Cm)^{Cm}}{(6n)^{m}}$
for all $m\geq 2k$ and $6n\geq Cm^{C}$.
@@
Let $C_{0}>1$ be the constant in Proposition 3.4. Assume $m\geq 2k$ and $6n\geq C_{0}m^{C_{0}}$. From (3) we know that
$\mathbb{E}_{n}\left[\textnormal{fix}_{\omega}\right]=\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\mathbb{E}_{n}(N_{\Gamma})$
. Then we have
@@
for the case that $\eta(\Gamma)\geq m$. Now apply Proposition 3.4 to $(g,\Gamma)\in\mathcal{I}(\omega)$ and $l=m-\eta(\Gamma)$, we have
(9) $\left|\mathbb{E}_{n}(N_{\Gamma})-\frac{1}{(6n)^{\eta(\Gamma)}}\sum\limits_{i=0}^{m-1-\eta(\Gamma)}\frac{a_{i}(\Gamma)}{(6n)^{i}}\right|\leq\frac{(C_{0}m)^{C_{0}(m-\eta(\Gamma))}}{(6n)^{m}}.$
@@
$\{(g,\Gamma)\in\mathcal{I}(\omega);\ \eta(\Gamma)=h\}=|\mathfrak{F}_{h+1}(\omega)|\leq k^{2h+2}.$
It then follows that for $m\geq 2k$ and $6n\geq(2C_{0}m)^{2C_{0}}$,
(10) $\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\frac{(C_{0}m)^{C_{0}(m-\eta(\Gamma))}}{(6n)^{m}}$ $=\sum\limits_{h=0}^{\infty}\sum\limits\limits_{\begin{subarray}{c}(g,\Gamma)\in\mathcal{I}(\omega)\\ \eta(\Gamma)=h\end{subarray}}\frac{(C_{0}m)^{C_{0}(m-h)}}{(6n)^{m}}$
$\leq\sum\limits_{h=0}^{\infty}\frac{(C_{0}m)^{C_{0}(m-h)}}{(6n)^{m}}\times k^{2h+2}$
$\leq\frac{(2C_{0}m)^{2C_{0}m}}{(6n)^{m}}.$
@@
where in the summation above, $a_{i}(\Gamma)=0$ for $i<0$. Together with (LABEL:e-tr-1), (9) and (10), one may conclude that $C=2C_{0}$ and $\{v_{j}(\gamma)\}_{j\geq 0}$ are desired. The proof is complete. ∎
@@ 4.2. Estimate of traces
There exist a universal constant $C\geq 1$ and a sequence $\{u_{i}(\gamma)\}_{i\geq 0}$ of real numbers for all $\gamma\in\textnormal{PSL}(2,\mathbb{Z})$ such that $u_{0}(\gamma)=\rho(\lambda(\gamma))$ and
$\left|\mathbb{E}_{n}\left[\textnormal{tr }\Pi_{\gamma}\right]-\sum\limits_{i=0}^{m-1}\frac{u_{i}(\gamma)}{(6n)^{i}}\right|\leq\frac{(Cm)^{Cm}}{(6n)^{m}}$
for all $m\geq|\gamma|+1$ and $6n\geq Cm^{C}$.
@@
Now we prove it case by case.
For the first case, i.e. $\gamma=e$, we have for any $(\sigma,\tau)\in\mathcal{E}_{n}$,
@@
Let $C$ be the universal constant in Proposition 3.7. From (12) and Proposition 3.7, we have for any $(m-1)\geq|\gamma|\geq 2k$ and $6n\geq Cm^{C}$,
@@
$=\frac{1}{6n}\left|\mathbb{E}_{n}[\textnormal{fix}_{\omega}]-\sum\limits_{i=0}^{m-2}\frac{v_{i}(\omega)}{(6n)^{i}}\right|\leq\frac{(Cm)^{Cm}}{(6n)^{m}}.$
@@ 4.3. Estimate of operator norms
Assume $x\in\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$. For $(\sigma,\tau)\in\mathcal{E}_{n}$, as $n\to\infty$
@@
in probability. The norm on each side is the operator norm.
Proof.
We follow the strategy in the proof of [MPvH25, Theorem 1.1]. From Proposition 4.1 and Proposition 4.3, an identical proof of [MPvH25, Theorem 1.5] gives that
@@
as $n\to\infty$ for any $x\in\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$. Next, we prove the lower bound. In [BCdlH94], it is proved that the reduced $C^{\star}$-algebra $\mathbb{C}^{\star}_{\textnormal{red}}[\textnormal{PSL}(2,\mathbb{Z})]$ is simple. Combining this fact and [BKKO17, Theorem 1.6], one may conclude that $\textnormal{PSL}(2,\mathbb{Z})$ has the unique trace property (one may also refer to [dlH00, Page 39]). It then follows from (18) and [MdlS24, Section 5.3] that
@@
The proof is complete. ∎
As a direct corollary, we have (one may also see e.g. [LM25, Corollary 1.10]),
Corollary 4.7.
For any $r\in\mathbb{N}$, finite supported map $a:\textnormal{PSL}(2,\mathbb{Z})\to M_{r\times r}(\mathbb{C})$ and random element $(\sigma,\tau)\in\mathcal{E}_{n}$,
@@
as $n\to\infty.$
The proof of Corollary 4.7 from Theorem 4.6 is similar to that in [HTr05, Section 9], which is based on the fact that there is a unique $C^{\star}$-algebra norm on $M_{r\times r}(\mathbb{C})$.
Finally, we have the following theorem.
Theorem 4.8 (=Theorem 1.1).
For any $\epsilon>0$,
$\lim\limits_{n\to\infty}\textnormal{Prob}^{n}_{\textnormal{BM}}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \lambda_{1}(S^{C}(\Gamma,\mathcal{O}))\geq\frac{1}{4}-\epsilon\right)=1.$
Proof.
Following [HM23], from Corollary 4.7 we have that for any $\epsilon>0$,
$\lim\limits_{n\to\infty}\textnormal{Prob}_{n}\left((\sigma,\tau)\in\mathcal{E}_{n};\ \begin{matrix}\textnormal{Spec}(\Delta_{\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z})})\cap\left(0,\frac{1}{4}-\epsilon\right)\vskip 5.0pt\\ =\textnormal{Spec}(\Delta_{S^{O}(\sigma,\tau)})\cap\left(0,\frac{1}{4}-\epsilon\right)\end{matrix}\right)=1.$
This together with Proposition 2.4 implies that
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \begin{matrix}\textnormal{Spec}(\Delta_{\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z})})\cap\left(0,\frac{1}{4}-\epsilon\right)\vskip 5.0pt\\ =\textnormal{Spec}(\Delta_{S^{O}(\Gamma,\mathcal{O})})\cap\left(0,\frac{1}{4}-\epsilon\right)\end{matrix}\right)=1$
Since (see e.g. [Ber11, Theorem 3.38])
$\lambda_{1}(\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z}))>\frac{1}{4},$
it follows that
(19) $\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \lambda_{1}\left(S^{O}(\Gamma,\mathcal{O})\right)>\frac{1}{4}-\epsilon\right)=1.$
From Proposition 2.1, Lemma 2.2 and (19), we have
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\sigma,\tau)\in\mathcal{F}_{n}^{\star};\ \lambda_{1}\left(S^{C}(\Gamma,\mathcal{O})\right)>\frac{1}{4}-\epsilon\right)=1.$
The proof is complete. ∎

### v2

@@ 2.3. An alternative description of Brooks-Makover model
For $n\geq 1$, assume that $S_{6n}$ is the permutation group of $6n$ elements. Now we consider a special subset of $S_{6n}\times S_{6n}$, which is defined as
@@
    of $T_{k}$. Then glue the sides $j_{s}^{1}$ and $j_{s}^{2}\ (1\leq s\leq 3n)$ together subject to two precise conditions;
@@
for any $\gamma\in\textnormal{PSL}(2,\mathbb{Z}),\ x\in\mathbb{H}$ and $i\in[6n]$. A Belyi surface can be viewed as a covering surface of the modular space (see [BM01, Section 1]) and [LP24, Subsection 2.6]. We provide a detailed proof in our setting here.
@@ 3.2.1. Injective $X$-labeled graph homomorphisms
The following lemma is the same as [MPvH25, Lemma 2.2] except an explicit expression for the constant $C^{\prime}$ in [MPvH25, Lemma 2.2].
@@
Let P,QP,\ Q be real polynomials of degree at most $m\in\mathbb{N}$, and let
@@
for all $n\in\mathbb{N}$ with $n\geq(Cm)^{C}$. Denote by $C^{\prime}=4C+5$, then for all $l\leq Cm$,
@@
Proof.
From the proof of [MPvH25, Lemma 2.2], we have
$\left|\Phi^{(l)}(t)\right|\leq 2^{l}(2C)^{2l+2}(2Cm)^{4Cl}l!$
for any $t\in[0,(2Cm)^{-2C-1}]$. Hence for any $l\leq Cm$,
$\sup\limits_{t\in\left[0,(2Cm)^{-2C-1}\right]}\left|\Phi^{(l)}(t)\right|$ $\leq 2^{l}(2C)^{2l+2}(2Cm)^{4Cl}l!$
$\leq(2C)^{2l+2}(2Cm)^{(4C+1)l}$
$\leq(2Cm)^{(4C+5)l}$
which implies that
$\sup\limits_{t\in\left[0,(C^{\prime}m)^{-C^{\prime}}\right]}\left|\Phi^{(l)}(t)\right|\leq\sup\limits_{t\in\left[0,(2Cm)^{-2C-1}\right]}\left|\Phi^{(l)}(t)\right|\leq(C^{\prime}m)^{C^{\prime}l}.$
The proof is complete. ∎
@@
    For all $m\geq 2k$, $l\leq m$ and $6n\geq(17m)^{17}$,
    $\left|\mathbb{E}_{n}(N_{\Gamma})-\frac{1}{(6n)^{\eta(\Gamma)}}\sum\limits_{i=0}^{l-1}\frac{a_{i}(\Gamma)}{(6n)^{i}}\right|\leq\frac{(17m)^{17l}}{l!(6n)^{l+\eta(\Gamma)}}.$
@@
(1). Assume that $\Gamma$ has $p$ $x_{1}$-cycles and $q$ $x_{2}$-cycles. Since $(g,\Gamma)\in\mathcal{I}(\omega)$, we have
@@
(2). Assume $m\geq 2k\geq 2p$. For $6n\geq(3m)^{3}$, we have
$\left|P\left(\frac{1}{6n}\right)\right|<3$
@@
$3$ $>Q\left(\frac{1}{6n}\right)=\prod\limits_{i=0}^{2p-1}\left(1-\frac{i}{6n}\right)$
$\geq\prod\limits_{i=0}^{2p-1}\left(1-\frac{i}{(3m)^{3}}\right)\geq\prod\limits_{i=0}^{2p-1}\left(1-\frac{1}{i^{2}}\right)>\frac{1}{3}.$
Apply Lemma 3.3 to P(t),Q(t)P(t),\ Q(t) and $C=3$, we have for any $l\leq 3m$,
$\sup\limits_{t\in\left[0,(17m)^{-17}\right]}\left|\left(\frac{P}{Q}\right)^{(l)}(t)\right|\leq(17m)^{17l}.$
It follows that for $6n\geq(17m)^{17}$ and $l\leq m$, there exists $\xi\in[0,\frac{1}{6n}]$ such that
@@ 3.2.2. Surjective $X$-labeled graph homomorphisms
Recall from the algorithm above the initial data $\mathcal{A}_{1}=(\Gamma_{1},G_{1},f_{1},a_{1})$, and after the $t$-th step ($1\leq t\leq k$), we obtain
@@
Moreover, there exists some $1\leq t_{0}\leq k$ such that at the $t_{0}$-th step, we make Operation II and join $w_{1}$ with some $w_{i}$ by an $x_{1}$-cycle. Since there are at most $\binom{k}{h}$ different choices to determine the remaining $h$ steps at which we make Operation II, it follows that
$|\mathfrak{F}_{h}(\omega)|\leq\binom{k}{h}\times k^{h-1}\leq k^{2h-1},$
@@ 3.2.3. Number of fixed points
where $k\geq 1$ and $i_{j}\in\{1,2\}\ (1\leq j\leq k)$. Then there exist a sequence $\{v_{i}(\omega)\}_{i\geq 0}\subset\mathbb{R}$ such that
$\left|\mathbb{E}_{n}\left[\textnormal{fix}_{\omega}\right]-\sum\limits_{i=0}^{m-1}\frac{v_{i}(\omega)}{(6n)^{i}}\right|\leq\frac{(18m)^{18m}}{(6n)^{m}}$
for all $m\geq 2k$ and $6n\geq 18m^{18}$.
@@
Assume $m\geq 2k$ and $6n\geq 17m^{17}$. From (5) we know that
$\mathbb{E}_{n}\left[\textnormal{fix}_{\omega}\right]=\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\mathbb{E}_{n}(N_{\Gamma}).$
Then we have
@@
for the case $\eta(\Gamma)\geq m$. Now applying Proposition 3.4 to $(g,\Gamma)\in\mathcal{I}(\omega)$ and $l=m-\eta(\Gamma)$, we have
(11) $\left|\mathbb{E}_{n}(N_{\Gamma})-\frac{1}{(6n)^{\eta(\Gamma)}}\sum\limits_{i=0}^{m-1-\eta(\Gamma)}\frac{a_{i}(\Gamma)}{(6n)^{i}}\right|\leq\frac{(17m)^{17(m-\eta(\Gamma))}}{(6n)^{m}}.$
@@
$\{(g,\Gamma)\in\mathcal{I}(\omega);\ \eta(\Gamma)=h\}=|\mathfrak{F}_{h+1}(\omega)|\leq k^{2h+1}.$
It then follows that for $m\geq 2k$ and $6n\geq 18m^{18}$,
(12) $\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\frac{(17m)^{17(m-\eta(\Gamma))}}{(6n)^{m}}$ $=\sum\limits_{h=0}^{\infty}\sum\limits\limits_{\begin{subarray}{c}(g,\Gamma)\in\mathcal{I}(\omega)\\ \eta(\Gamma)=h\end{subarray}}\frac{(17m)^{17(m-h)}}{(6n)^{m}}$
$\leq\sum\limits_{h=0}^{\infty}\frac{(17m)^{17(m-h)}}{(6n)^{m}}\times k^{2h+1}$
$\leq\frac{(18m)^{18m}}{(6n)^{m}}.$
@@
where in the summation above, $a_{i}(\Gamma)=0$ for $i<0$. Together with (LABEL:e-tr-1), (11) and (12), one may conclude that $\{v_{j}(\gamma)\}_{j\geq 0}$ is as desired. The proof is complete. ∎
@@
Remark.
We thank Doron Puder for informing us that he and Zimhoni, in their work [PZ24], studied a more general model involving two independent random permutations $\sigma$ and $\tau$, where $\sigma^{2}=\textnormal{id}$ and $\tau^{3}=\textnormal{id}$, both of which are allowed to have nonempty fixed points. Although the overall behavior of these two models differs significantly, certain similarities remain. Interested readers may refer to [PZ24] for more details.
@@ 4.2. Estimate of traces
There exists a sequence $\{u_{i}(\gamma)\}_{i\geq 0}$ of real numbers for all $\gamma\in\textnormal{PSL}(2,\mathbb{Z})$ such that $u_{0}(\gamma)=\rho(\lambda(\gamma))$ and
$\left|\mathbb{E}_{n}\left[\textnormal{tr }\Pi_{\gamma}\right]-\sum\limits_{i=0}^{m-1}\frac{u_{i}(\gamma)}{(6n)^{i}}\right|\leq\frac{(18m)^{18m}}{(6n)^{m}}$
for all $m\geq|\gamma|+1$ and $6n\geq 18m^{18}$.
@@
Now we prove it case by case. For the first case, i.e. $\gamma=e$, we have for any $(\sigma,\tau)\in\mathcal{E}_{n}$,
@@
From (14) and Proposition 3.7, we have for any $(m-1)\geq|\gamma|\geq 2k$ and $6n\geq 18m^{18}$,
@@
$=\frac{1}{6n}\left|\mathbb{E}_{n}[\textnormal{fix}_{\omega}]-\sum\limits_{i=0}^{m-2}\frac{v_{i}(\omega)}{(6n)^{i}}\right|\leq\frac{(18m)^{18m}}{(6n)^{m}}.$
@@ 4.3. Estimate of operator norms
In this subsection, we complete the proof of Theorem 1.1. We first prove the strong convergence property of $\textnormal{PSL}(2,\mathbb{Z})$.
Recall that for any $(\sigma,\tau)\in\mathcal{E}_{n}$
$\Pi(\sigma,\tau)=\textnormal{std}\circ\Phi(\sigma,\tau):\textnormal{PSL}(2,\mathbb{Z})\to\mathcal{U}(6n-1),$
and it could be extended to $\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$ naturally. For any real polynomial $h$ and $x\in\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$, $h\left(\Pi(\sigma,\tau)(x)\right)$ is a matrix. Denote by
$\textnormal{tr }h\left(\Pi(\sigma,\tau)(x)\right)=\frac{1}{6n}\textnormal{Tr }h\left(\Pi(\sigma,\tau)(x)\right).$
Then $\textnormal{tr }h(\Pi_{\gamma}):\mathcal{E}_{n}\to\mathbb{R}$ defined by
$\textnormal{tr }h(\Pi_{x})(\sigma,\tau)=\frac{1}{6n}\textnormal{Tr }h\left(\Pi(\sigma,\tau)(x)\right)$
is a random variable on $\mathcal{E}_{n}$. From Proposition 4.1 and an identical proof of [MPvH25, Lemma 3.1], we have the following property.
Proposition 4.6.
[…]
and
[…]
for all $n\geq 1$ and every real polynomial $h$ of degree at most $q$.
Remark.
From (5) and (7) we know that for any $\omega\in\mathbb{F}_{2}$,
$\mathbb{E}_{n}[\textnormal{fix}_{\omega}]=\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\mathbb{E}_{n}(N_{\Gamma})=\sum\limits_{(g,\Gamma)\in\mathcal{I}(\omega)}\frac{1}{(6n)^{\eta(\Gamma)}}\times\frac{\prod\limits_{i=0}^{p-1}\left(1-\frac{2i}{6n}\right)\times\prod\limits_{i=1}^{q-1}\left(1-\frac{3i}{6n}\right)}{\prod\limits_{i=0}^{2p-1}\left(1-\frac{i}{6n}\right)}$
is a rational function of $\frac{1}{6n}$. One may also deduce Proposition 4.6 by directly applying the methods in [CGVTvH24] without using Proposition 4.1. In this paper, the important function $u_{1}$ in Proposition 4.3, which is essential in the proof of Theorem 4.8, is defined in Proposition 4.1. We are grateful to Will Hide, Doron Puder and Ramon van Handel for kindly pointing this to us.
From Proposition 4.6, the proof of [MPvH25, Corollary 3.2] gives the following result.
Corollary 4.7.
[…]
for all $n\geq 1$, $h\in C^{\infty}(\mathbb{R})$ and $f(\theta)=h(K\cos\theta)$.
@@
For any self-adjoint $x\in\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$, $n\geq 1$ and $\epsilon>0$,
$\mathrm{Prob}^{n}_{\mathrm{BM}}\Big((\sigma,\tau)\in\mathcal{E}_{n};\ \big\|\Pi(\sigma,\tau)(x)\big\|\geq(1+\epsilon)\big\|\lambda(x)\big\|\Big)\leq\frac{C(x)}{n\epsilon^{217}}$
where $C(x)>0$ is a constant that only depends on $x$.
Proof.
@@
and
(21) $f^{(m+1)}||_{[0,2\pi]}\leq 16(4m)^{m}\left(\frac{2K}{\delta}\right)^{m+\frac{1}{2}}$.
From Proposition 4.3 and [CGVTvH24, Lemma 4.9], we have
@@
which together with (20), (21) and Corollary 4.7 implies that
$\rho(h(\lambda(x)))=\nu(h)=0,$
and
$\ \ \ \ \mathrm{Prob}^{n}_{\textnormal{BM}}\left((\sigma,\tau)\in\mathcal{E}_{n};\ \big\|\Pi(\sigma,\tau)(x)\big\|\geq\big\|\lambda(x)\big\|+\delta\right)$
$\leq\mathbb{E}_{\textnormal{BM}}^{n}\left[\textnormal{Tr\ }h(\Pi(\sigma,\tau))(x)\right]$
$\leq 6n\times\frac{32(108|x|)^{216}}{(6n)^{2}}\times 16\times 864^{216}\left(\frac{2K}{\delta}\right)^{216+\frac{1}{2}}$
$=O\left(\frac{1}{n}\left(\frac{|x|K}{\delta}\right)^{217}\right).$
Apply [MPvH25, Equation (3.3)] for $h(t)=t$ and $q=1$, we have
@@
where $x=\sum\limits_{\gamma}a_{\gamma}\cdot\gamma$. It follows that
$\mathrm{Prob}^{n}_{\textnormal{BM}}\left((\sigma,\tau)\in\mathcal{E}_{n};\ \big\|\Pi(\sigma,\tau)(x)\big\|\geq(1+\epsilon)\big\|\lambda(x)\big\|\right)\leq\frac{C(x)}{n\epsilon^{217}},$
where $C(x)$ is a constant that only depends on $x$. The proof is complete. ∎
Based on Theorem 4.8, the proof of [MPvH25, Theorem 3.4] gives the following theorem, which is a matrix coefficient version of Theorem 4.8.
Theorem 4.9.
For any self-adjoint $x\in M_{d}(\mathbb{C})\otimes\mathbb{C}[\textnormal{PSL}(2,\mathbb{Z})]$, $n\geq 1$ and $\epsilon>0$, we have
[…]
Finally, we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
For $r>0$, set
$\lambda(r,n)=\frac{1}{\sqrt{r}}+\left(\frac{r\log n}{n}\right)^{\frac{1}{217}}.$
By Theorem 2.8 and Theorem 4.9, the proof of [HMT25a, Theorem 1.1] gives that there exists a universal constant $c>0$ such that
(22) $\lim\limits_{n\to\infty}\textnormal{Prob}_{n}\left((\sigma,\tau)\in\mathcal{E}_{n};\ \begin{matrix}\textnormal{Spec}(\Delta_{\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z})})\cap\left(0,\frac{1}{4}-c\lambda(r,n)\right)\vskip 5.0pt\\ =\textnormal{Spec}(\Delta_{S^{O}(\sigma,\tau)})\cap\left(0,\frac{1}{4}-c\lambda(r,n)\right)\end{matrix}\right)=1.$
Now we take $r=n^{\frac{1}{110}}$. Then from Proposition 2.7 and (22) we have
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \begin{matrix}\textnormal{Spec}(\Delta_{\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z})})\cap\left(0,\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)\vskip 5.0pt\\ =\textnormal{Spec}(\Delta_{S^{O}(\Gamma,\mathcal{O})})\cap\left(0,\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)\end{matrix}\right)=1.$
Since $\lambda_{1}(\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z}))>\frac{1}{4}$ (see e.g. [Ber11, Theorem 3.38]), it follows that
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \textnormal{Spec}(\Delta_{S^{O}(\Gamma,\mathcal{O})})\cap\left(0,\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)=\emptyset\right)=1.$
This together with Proposition 2.1 completes the proof. ∎